OpenStudy (anonymous):

divide and simplify 10 over x-5 minus 10 over 5-x

OpenStudy (anonymous):

10/(x-5)-10/(5-x) =10/(x-5)-(-10/(x-5)) (factoring out a negative one from the bottom and applying it to the top. This is the same as saying (-2)/3=2/(-3) =10/(x-5)+10/(x-5) =20/(x-5)